The election returns showed that a certain candidate received 65% of the 

votes. Find the probability that two random samples each consist of 200 

voters, indicated a greater than 10% difference in the proportions that voted 

for the candidate.



1
Expert's answer
2021-06-24T16:19:01-0400

P=0.65

N1=n2=200

Standard deviation-=√{P(1-p)/n1+ p(1-p)/n2}

   =√{(0.65*0.35)/200*2}= 0.0477

Z={(p1-p2)-(P1-P2)}/standard deviation

 But p1-p2= 0.1 and P1-P2=0.65-0.65=0

Z=0.1/0.0477=2.096

P(Z>2.096)= 1-0.9821

=0.01179